OpenStudy (anonymous):

What is the slope for the following table of values? The table represents the total amount spent in a paperback book club after an introductory fee of $12. # books purchased 1 2 3 4 Cost of book club 21 30 39 48 A. 3 B. 5 C. 7 D. 9

OpenStudy (anonymous):

9

OpenStudy (anonymous):

its very simple apply ur mind with this 1, What is the slope for the following table of values? gallons of gas used 1 2 3 4 miles driven 26 52 78 104 A, 2 B, 4 C, 30 D, 26 2, Write an equation in y = mx + b form for the table. gallons of gas used 1 2 3 4 miles driven 26 52 78 104 A, y = 26x B, y = 26x + 26 C, y = 26x + 1 D, y = 26x -1 3, Use the equation for the table to predict the number of miles you could drive on 12 gallons of gas. gallons of gas used 1 2 3 4 miles driven 26 52 78 104 A, 200 B, 252 C, 300 D, 312 4, What is the slope for the following table of values? The table represents the total amount spent in a paperback book club after an introductory fee of $12. Books bought 1 3 5 7 Total Amount Spent $ 21 39 57 75 A, 3 B, 5 C, 7 D, 9 5, Write an equation in y = mx + b form for the table. Books Bought 1 3 5 7 Total Amount Spent $ 21 39 57 75 A, y = 3x + 21 B, y = 9x + 12 C, y = 5x + 21 D, y = 7x + 12 1 year ago Report Abuse 1) D, 26 2) A, y = 26x 3) D, 312 4) D, 9 5) B, y = 9x + 12

OpenStudy (anonymous):

understood its 9